OpenStudy (anonymous):

Sofia joined a golf club two years ago. She pays an annual membership fee of $1,045 and a greens fee of $35 each time she plays a game of golf. The function below can be used to calculate the total yearly golfing fee, f (g), in dollars. f (g) = 1045 + 35g, where g represents the number of times she played golf during the year. Last year she paid $1,710 as a total golfing fee. For how many games did she pay a greens fee?

OpenStudy (anonymous):

She paid 19 greens fee

OpenStudy (luigi0210):

no work?

OpenStudy (anonymous):

fffunction(fee) = total yearly fee + greens fee(number of times she played golf) f(g) = 1045 + 35g f(1,710) = 1045 + 35g Isolate 35g. 665 = 35g g = 19 She paid 19 games for a greens fee. understand?

OpenStudy (luigi0210):

Another person enlighten

OpenStudy (anonymous):

Thanks :D

OpenStudy (luigi0210):

Good job Miss Callie ^_^

OpenStudy (anonymous):

XD lol thanx